Three brothers were born 3 years in a row. The sum of their ages is 57.How old is the youngest brother?
Answer by macston(5194)

.
Y=age of youngest; M=age of middle=y+1; O=age of oldest=Y+2
Y+M+O=57
Y+(Y+1)+(Y+2)=57
3Y+3=57
3Y=54
Y=18
.
ANSWER: The youngest brother is 18 years old.
.
CHECK:
Y=18
M=Y+1=18+1=19
O=y+2=18+2=20
.
Y+M+O=57
18+19+20=57
57=57